OpenStudy (anonymous):

What is the value of this expression when c = 6? 3c - 4 / 2 A. 14 B. 11 C. 7 D. 5 I did 3c - 4 / 2 = 3(6) - 4 / 2 = 18 - 4 / 2 = 14 / 2 = 7. Am I correct?

Directrix (directrix):

 I did 3*6 - 4/2 = 18 - 2 = 16 because the problem implies 3c - (4/2). You did 3c - 4 / 2 = 3(6) - 4 / 2 = 18 - 4 / 2 = 14 / 2 = 7 because you thought the problem implied [3c - 4] / 2. Grouping symbols are essential for clarity in mathematics.